Find the change in entropy when (a) 3.0 kg of boiling water changes to steam at 100ºC, (b) 3.0 kg of water is heated from 20ºC to 80ºC, and (c) 2.0 kg of water at 30ºC are mixed with 5.0 kg of water at 75ºC and the mixture allowed to reach equilibrium.

1
Expert's answer
2022-04-26T15:19:20-0400

a)

"S=\frac{Lm}T=18.5~\frac{kJ}K,"

b)

"S=cm=12.6~\frac{kJ}K,"

c)

"S=\frac 0T=0~J."